# Hierarchies of direct powers, ultrapowers and cumulative powers

### 16 Mar 2026 v1

@@ Abstract.
In this paper we investigate cumulative hierarchies of functions on structures, or cumulative powers, and study their properties. Particularly, we show how they extend the preservation phenomena of reduced powers, direct powers and ultrapowers by offering a characterization result of the fragment of first-order theory it preserves, and elucidate the connection between the three sorts of constructions. More precisely, we show how both direct powers and ultrapowers may be obtained from cumulative powers as a quotient by appropriate equivalence relations. We address how embeddability lifts from generating structures to their cumulative powers, direct powers and ultrapowers, and under what conditions ultrapowers embed into corresponding cumulative powers or direct powers. We further offer an application of the framework to show a straightforward way of constructing a surreal field.
@@ 1. Introduction
Consider the collection of all functions from a given index set $U$ into the naturals, $\mathbb{N}^{I}$. There is a straightforward manner of extending any operation $\mathrm{F}$ on $\mathbb{N}$ to a $\mathrm{F}^{\prime}$ on $\mathbb{N}^{I}$ that behaves similarly to the original operation, as in the corresponding direct power – that is, pointwisely, so that $\mathrm{F}^{\prime}(f_{1},...,f_{n})=g$ iff for all $j\in I$, $\mathrm{F}\big(f_{1}(j),...,f_{n}(j)\big)=g(j)$. Likewise, we might define the constant function $\overline{0}$ as the identity element of the pointwise addition, and its image under the extended successor function is precisely the identity element of the extended multiplication. We may extrapolate this construction, and given a family of index sets $\{I_{n}\}_{n\in\omega}$, inductively defining $\mathbb{N}_{n}$ by setting $\mathbb{N}_{0}=\mathbb{N}$ and $\mathbb{N}_{n+1}=\mathbb{N}_{n}^{I_{n}}$, we may keep extending those operations and identity elements accordingly. Since there is a natural embedding between any two structures so defined, we may even take $\mathbb{N}_{\omega}$ to be the natural limit of embeddings, and given a proper class sized family of index sets $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$, keep transifinitely extending this hierarchy. Each of the $\mathbb{N}_{\beta}$ inherit nice properties of $\mathbb{N}$, and given the preservation results of direct powers by Alfred Horn [13], it may be checked that they are indeed semirings. We may do the same thing with $\mathbb{Z}$, and similarly get a hierarchy of rings $\mathbb{Z}_{\beta}$.
@@
Suppose now we take a cumulative counterpart of the above structures. Starting with $\mathbb{N}_{0}=\mathbb{N}$, we take $\mathbb{N}_{\beta^{+}}=\mathbb{N}_{\beta}^{I_{\beta}}\cup\mathbb{N}_{\beta}$. Then, the operations and constants may no longer be pointwisely defined, but given a definition according to which each stage of the hierarchy is semantically conservative with respect to its predecessors, for limit ordinals $\lambda$ we may take $\mathbb{N}_{\lambda}$ to be the simple union of all the preceding stages. Call such construction a cumulative (direct) power. Is there a natural way of extending addition and multiplication to those structures? If so, what first-order properties are retained by cumulative powers?
@@
The present paper has four different goals, interrelated as a broad study of cumulative powers and hierarchical constructions of functions. The first is to answer the above questions. To do so, in Section 2, we define a framework for working with hierarchies of cumulative powers of arbitrary structures. The main feature of the construction, offered in Definition 2.3, is a hereditary definition of the extended operations and relations of higher levels, so that they are reduced to the lower levels at which they are originally defined. The resulting structures have elements from different levels of such hierarchy, such as higher-order functions and non-functional elements, which nevertheless interact in a natural manner. We then show how a substantial fragment of first-order theory is preserved by the constructions. In fact, we offer the precise fragment of first-order theory preserved by cumulative powers. Call a sentence preserved by direct powers a direct power sentence. Let a formula be non-collapsible if its positive equalities do not force the collapse of different levels of the hierarchy of cumulative powers (as we show, a syntactic definition of non-collapsible formulas may be offered in Definition 2.17). Then:
@@
This extends well known results of preservation of first-order properties by reduced products by [4], [14], and [10], of direct products by [19], and the classical result for ultrapowers.
Given the construction of cumulative powers, the second aim of this paper naturally arises, which is how the new construction relates to direct powers and ultrapowers. As we argue in Section 3, the main difference between cumulative powers and direct powers is that in the latter equality is also defined hereditarily, whereas in the former, equality is defined as identity. This asymmetry ultimately accounts for the difference between the preservation results for direct powers and the one presented in Corollary 2.27. This fact is consolidated by showing, in Theorem 3.6, how quotienting a cumulative power of a structure by an equivalence relation, defined by an equality modulo hereditary identity, results in a structure which is isomorphic to its relative direct power.
In Section 4, we show a cumulative power may be further refined by quotienting it by yet another equivalence relation, given by an adequately defined ultrafilter, so that the resulting structure is isomorphic to a relative ultrapower, and therefore preserves any first-order property of a finite language. In this way, we may see how cumulative powers serve as a generalisation of both ultrapower and direct power constructions, as the choice of the equivalence relation which quotients it defines the first-order faithfulness to the original structure. The preservation results may be thus summarized as follows. Let $\mathcal{L}_{\mathcal{F}}$ be the fragment of first-order logic preserved by cumulative powers, $\mathcal{L}_{\Pi/F}$ the fragment preserved by reduced powers, $\mathcal{L}_{\Pi}$ the fragment preserved by direct powers, and $\mathcal{L}_{\Pi/\mathcal{U}}$ the fragment preserved by ultrapowers. Then, we have the following graph of containment of preserved first-order formulas:
@@
Cumulative powers, therefore, provide a structure from which we may take both direct powers and ultrapowers to arise by imposing appropriate equivalence relations, providing thus a natural hierarchy of preservation of first-order theory represented by the following flow chart. Let $I$ be an index set, $\mathcal{A}$ a structure, $\mathcal{F}_{1}(\mathcal{A})$ the cumulative power generated by them, and $\mathcal{U}$ an ultrafilter over $I$. Then:
@@
The third aim of the paper is to study how embeddability is lifted from structures to its cumulative powers, and resulting quotient structures. For natural reasons, part of that question is inevitably reduced to criteria of embedding between ultrapowers. [^1] Since the matter between direct powers is straightforward, as in Theorem 3.8. Let $\leq_{\mathrm{RK}}$ denote the Rudin-Keisler ordering of ultrafilters [17]. A partial answer is then offered by the following result:
@@
Given the equivalence results between quotiented cumulative powers and ultrapowers, with Corollary 4.22 gives us the desired result.
@@
As the last goal, we offer an application of the hierarchical constructions to real closed fields. We show that, at each ordinal stage of either a hierarchy generated by ultrapowers or quotiented cumulative powers, a suitable choice of ultrafilters produces hyperreal fields, whereas the hierarchy as whole, under the assumption of Global Choice – following results by Conway [5] and Ehrlich [7] –, forms a real closed field isomorphic to the field of surreal numbers.
@@ 1.1. Conventions
Before we begin, we start with a few remarks on conventions we shall use. For $n\in\omega$, a set $A$ and $a\in A$, we write $\langle x_{1},...,a,...,x_{n-1}\rangle$ for a $n$-tuple of $A^{n}$ in which $a$ occurs in some position. We shall also follow the convention of denoting structures by a calligraphic font, and its corresponding domain, by an italicized font – for example, $\mathcal{A}$ and $A$, respectively. Let $\mathcal{A}$ be a structure and $I$ be a set. We denote by $\mathcal{A}^{I}$ the direct power of $A$ with index set $I$, and by $A^{I}$ its domain, that is, the set of all functions from $I$ into $A$. If $\mathcal{U}$ is an ultrafilter over $I$, we denote by $\mathcal{A}^{I}/\mathcal{U}$ the resulting ultrapower. By the natural embedding of a set or structure into its direct power or ultrapower, we mean the mapping which takes each element of the domain of the set and maps it to either the constant function which assigns to every element of the index set the element of the domain in question, in the case of direct powers, or its corresponding equivalence class modulo the ultrafilter, in the case of ultrapowers. We let $\mathbf{On}$ stand for the class of ordinals, $\alpha$, $\beta$, $\delta$, $\lambda$ and so on, stand for ordinals, $\kappa$, $\tau$, $\xi$, $\zeta$, and so on, for cardinals, $\varphi$, $\psi$, $\gamma$ and so on, for formulas, $\Delta$ for an arbitrary string of quantifiers, and $t$ and $s$ for arbitrary terms of a given language. For an ordinal $\beta$ or a cardinal $\kappa$, we denote by $\beta^{+}$ and $\kappa^{+}$ their respective successors.
@@
We take the basic first-order language to be composed by individual variables, negation, conjunction, equality, and the universal quantifier (where the other connectives and quantifiers may be defined as usual). For an ordered set $\mathcal{A}$, we denote by $\mathrm{coi}(\mathcal{A})$ and $\mathrm{cof}(\mathcal{A})$ its coinitiality and cofinality, respectively. For a signature $\sigma$, we denote by $\mathcal{L}^{\sigma}$ the first-order language induced by it. As usual, we use a calligraphic letter $\mathcal{A}$ to denote a structure, and the respective italicized letter $A$ to denote its corresponding domain. Given a constant $\mathrm{c}$, function symbol $\mathrm{F}$ or relation symbol $\mathrm{R}$ of $\sigma$, and a $\sigma$-structure $\mathcal{A}$, we denote by $\mathrm{c}^{\mathcal{A}}$, $\mathrm{F}^{\mathcal{A}}$ and $\mathrm{R}^{\mathcal{A}}$ their interpretations in $\mathcal{A}$, respectively. Whenever we are working with some model, we shall state things like $\mathcal{A}\models\varphi[a_{1},...,a_{n}]$, for $a_{0},...,a_{n}\in A$, to mean $\varphi$ is a formula with at most $n$ free variables $x_{1},...,x_{n}$, and $\varphi$ is true in $\mathcal{A}$ when the $x_{i}$ are assigned the respective values $a_{i}$. Likewise, given a $\sigma$-term $t$, we shall denote by $t^{\mathcal{A}}(a_{1},...,a_{n})$ the element of $\mathcal{A}$ obtained by interpreting $t$ with each of its free variable $x_{i}$ being assigned the value $a_{i}$ – for example, in $\mathbb{N}$, $x[2]+y[3]$ is the element $2+^{\mathbb{N}}3=5$.
@@ 2. A hierarchy of cumulative powers
Let $\sigma=\langle\{F_{i}\}_{i\in I},\{R_{i}\}_{i\in J},\{c_{i}\}_{i\in K},\mathrm{ar}\rangle$ be a signature. [^4] Where $\mathrm{ar}:(\{F_{i}\}_{i\in I}\cup\{R_{i}\}_{i\in J})\to\mathbb{N}$ is the arity function. Let also $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of sets, $\mathcal{A}=\langle A,\langle\mathrm{F}_{i}\rangle_{i\in I},\langle\mathrm{R}_{i}\rangle_{i\in J},\langle\mathrm{c}_{i}\rangle_{i\in K}\rangle$ be a $\sigma$-structure, where $A$ is the underlying set, $\mathrm{F}_{i}:A^{\mathrm{ar}(F_{i})}\to A$ a $\mathrm{ar}(F_{i})$-nary function, $\mathrm{R}_{i}\subseteq A^{\mathrm{ar}(R_{i})}$ a $\mathrm{ar}(R_{i})$-ary relation, and $\mathrm{c}_{i}\in A$ the interpretation of the constant $c_{i}$. The cumulative power hierarchy generated by $\mathcal{A}$ is composed by the following structures, for each $\beta\in\mathbf{On}$:
@@
Notice the general scheme of the definition above: for a given atom $\varphi$ and elements $a_{1},...,a_{n}\in\mathcal{F}_{\beta}(A)$, $\varphi(a_{1},...,a_{n})$ is the case iff for any $j\in I_{\delta-1}$, $\varphi\big(\vartheta^{\delta}_{j}(a_{1}),...,\vartheta^{\delta}_{j}(a_{n})\big)$ holds, where $\delta=\linebreak\max\{\rho(a_{i})\}_{1\leq i\leq n}$. [^5] Although, in the case of the extended functions, one more constraint is imposed – and equality, of course, does not accord to that. Let us call such sort of definition a hereditary definition. We shall shortly see what may be said of structures defined hereditarily.
@@
    Notice that means the definition of $\mathrm{F}_{\beta}(a,b)$ when $\rho(a)\neq\rho(b)$ is such that $\mathrm{F}_{\beta}(a,b)=\mathrm{F}_{\gamma}(a,b)$, for every $\gamma\in[\delta,\beta]$. The intuitive idea in the definitions, represented by the function $\vartheta$, consists in going back to where the largest element is defined and use the operation there, which by assumption is already defined. In the definition of the relations, an equivalent definition would need to break down into the levels of each member of any $n$-tuple. For example, for a binary $\mathrm{R}$, an equivalent definition is:
@@
    We may see the amount of clauses in a definition like that grows proportionally to the number of permutations of a $n$-tuple, so that using the $\vartheta$ function allows for a more economic definition. Likewise, the definition of the set defined by the constants may be easily calculated to be:
@@
By induction on $\beta$. The base case is trivial, and for the inductive step it suffices to notice the definition of $\mathrm{F}_{\beta}$ works iff the $\mathrm{F}_{\alpha}$ are defined for all $\alpha<\beta$. The limit case is straightforward, as two elements in the limit case must be members of some iteration at which the operation is defined for them. ∎

### 18 Mar 2026 v2

@@ Abstract.
In this paper we investigate cumulative hierarchies of functions on structures, or cumulative powers, and study their properties. Particularly, we show how they extend the preservation phenomena of reduced powers, direct powers and ultrapowers by offering a characterization of the fragment of first-order theory it preserves, and elucidate the connections between the three sorts of constructions. More precisely, we show how both direct powers and ultrapowers may be obtained from cumulative powers as quotients by appropriate equivalence relations. We address how embeddability lifts from generating structures to their cumulative powers, direct powers and ultrapowers, and under what conditions ultrapowers embed into corresponding cumulative powers or direct powers. We further offer an application of the framework to show a straightforward way of constructing Conway’s surreal field.
@@ 1. Introduction
Consider the collection of all functions from a given index set $I$ into the naturals, $\mathbb{N}^{I}$. There is a straightforward manner of extending any operation $\mathrm{F}$ on $\mathbb{N}$ to an $\mathrm{F}^{\prime}$ on $\mathbb{N}^{I}$ that behaves similarly to the original operation, as in the corresponding direct power – that is, pointwisely, so that $\mathrm{F}^{\prime}(f_{1},...,f_{n})=g$ iff for all $j\in I$, $\mathrm{F}\big(f_{1}(j),...,f_{n}(j)\big)=g(j)$. Likewise, we might define the constant function $\overline{0}$ as the identity element of the pointwise addition, and its image under the extended successor function is precisely the identity element of the extended multiplication. We may extrapolate this construction, and given a family of index sets $\{I_{n}\}_{n\in\omega}$, by inductively defining $\mathbb{N}_{n}$ by setting $\mathbb{N}_{0}=\mathbb{N}$ and $\mathbb{N}_{n+1}=(\mathbb{N}_{n})^{I_{n}}$, we may keep extending those operations and identity elements accordingly. Since there is a natural embedding between any two structures defined in this way, we may even take $\mathbb{N}_{\omega}$ to be the direct limit of embeddings, and given a proper class sized family of index sets $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$, keep transfinitely extending this hierarchy. Each of the $\mathbb{N}_{\beta}$ inherits nice properties of $\mathbb{N}$, and given the preservation results of direct powers by Alfred Horn [12] and Joseph Weinstein [20], it may be checked that they are indeed semirings. We may do the same thing with $\mathbb{Z}$, and similarly get a hierarchy of rings $\mathbb{Z}_{\beta}$.
@@
Suppose now we take a cumulative counterpart of the above structures. Starting with $\mathbb{N}_{0}=\mathbb{N}$, we take $\mathbb{N}_{\beta^{+}}=(\mathbb{N}_{\beta})^{I_{\beta}}\cup\mathbb{N}_{\beta}$. Then, the operations and constants may no longer be pointwisely defined, but given a definition according to which each stage of the hierarchy is semantically conservative with respect to its predecessors, for limit ordinals $\lambda$ we may take $\mathbb{N}_{\lambda}$ to be the simple union of all the preceding stages. Call such a construction a cumulative (direct) power. Is there a natural way of extending addition and multiplication to those structures? If so, what first-order properties are retained by cumulative powers?
@@
The present paper has four different goals, interrelated within a broad study of cumulative powers and hierarchical constructions of functions. The first is to answer the above questions. To do so, in Section 2, we define a framework for working with hierarchies of cumulative powers of arbitrary structures. The main feature of the construction, offered in Definition 2.3, is a hereditary definition of the extended operations and relations of higher levels, so that they are reduced to the lower levels at which they are originally defined. The resulting structures have elements from different levels of such a hierarchy, such as higher-order functions and non-functional elements, which nevertheless interact in a natural manner. We then show how a substantial fragment of first-order theory is preserved by the constructions by offering the precise fragment of first-order theory preserved by cumulative powers. To present it, let us call a sentence preserved by direct powers a direct power sentence (a characterization of which is offered in [20]). Let a formula be non-collapsible if its positive equalities do not force the collapse of different levels of the hierarchy of cumulative powers – a property which, as given in Definition 2.18, may be syntactically defined. Then:
@@
This extends the well known results of preservation of first-order properties by reduced products by [4], [13], and [9], of direct products by [20], and the classical Łoś’ theorem for ultraproducts [16]. These preservation results may be thus summarized as follows. Let $\mathcal{L}_{\mathcal{F}}$ be the fragment of first-order logic preserved by cumulative powers, $\mathcal{L}_{\Pi/F}$ the fragment preserved by reduced powers, $\mathcal{L}_{\Pi}$ the fragment preserved by direct powers, and $\mathcal{L}_{\Pi/\mathcal{U}}$ the fragment preserved by ultrapowers. Then, we have the following graph of containment of preserved first-order formulas:
@@
Given the construction of cumulative powers, the second aim of this paper naturally arises, which is how the new construction relates to direct powers and ultrapowers. As we argue in Section 3, the main difference between cumulative powers and direct powers is that in the latter equality is also defined hereditarily, whereas in the former, equality is defined as identity. This asymmetry ultimately accounts for the difference between the preservation results for direct powers and the one presented in Corollary 2.27. This fact is consolidated by showing, in Theorem 3.6, how taking a quotient of a cumulative power of a structure by an equivalence relation, defined by an equality modulo hereditary identity, results in a structure which is isomorphic to its relative direct power.
In Section 4, we show a cumulative power may be further refined by taking a quotient of it by yet another equivalence relation, given by an adequately defined ultrafilter, so that the resulting structure is isomorphic to a relative ultrapower, and therefore preserves any first-order property of a finite language. In this way, we may see how cumulative powers serve as a generalisation of both ultrapower and direct power constructions, as the choice of the equivalence relation which quotients it defines the first-order faithfulness to the original structure. We may appreciate the connection between the different structures in the following way. Let $I$ be an index set, $\mathcal{A}$ a structure, $\mathcal{F}_{1}(\mathcal{A})$ the cumulative power generated by them, and $\mathcal{U}$ an ultrafilter over $I$. We may thus see that:
@@
Cumulative powers, therefore, provide a structure from which we may take both direct powers and ultrapowers to arise by imposing appropriate equivalence relations.
The third aim of the paper is to study how embeddability is lifted from structures to their cumulative powers, and resulting quotient structures. For natural reasons, part of that question is inevitably reduced to criteria of embedding between ultrapowers. [^1] Since the matter between direct powers is straightforward, as in Theorem 3.8. Let $\leq_{\mathrm{RK}}$ denote the Rudin-Keisler ordering of ultrafilters [17]. A partial answer is then offered by the following result:
@@
Given the equivalence results between quotiented cumulative powers and ultrapowers, Corollary 4.22 gives us the desired result.
@@
As the last goal, we offer an application of the hierarchical constructions to real closed fields. We show that, at each ordinal stage of either a hierarchy generated by ultrapowers or quotiented cumulative powers, a suitable choice of ultrafilters produces hyperreal fields, whereas the hierarchy as a whole, under the assumption of Global Choice – following results by Conway [5] and Ehrlich [6] –, forms a real closed field isomorphic to the surreal numbers.
@@ 1.1. Conventions
Before we begin, we start with a few remarks on conventions we shall use. For $n\in\omega$, a set $A$ and $a\in A$, we write $\langle x_{1},...,a,...,x_{n-1}\rangle$ for an $n$-tuple of $A^{n}$ in which $a$ occurs in some position. We shall also follow the convention of denoting structures by a calligraphic font, and their corresponding domains, by an italicized font – for example, $\mathcal{A}$ and $A$, respectively. Let $\mathcal{A}$ be a structure and $I$ be a set. We denote by $\mathcal{A}^{I}$ the direct power of $A$ with index set $I$, and by $A^{I}$ its domain, that is, the set of all functions from $I$ into $A$. If $\mathcal{U}$ is an ultrafilter over $I$, we denote by $\mathcal{A}^{I}/\mathcal{U}$ the resulting ultrapower. By the natural embedding of a set or structure into its direct power or ultrapower, we mean the mapping which takes each element of the domain of the set and maps it to either the constant function which assigns to every element of the index set the element of the domain in question, in the case of direct powers, or its corresponding equivalence class modulo the ultrafilter, in the case of ultrapowers. We let $\mathbf{On}$ stand for the class of ordinals, $\alpha$, $\beta$, $\delta$, $\lambda$ and so on, stand for ordinals, $\kappa$, $\tau$, $\xi$, $\zeta$, and so on, for cardinals, $\varphi$, $\psi$, $\gamma$ and so on, for formulas, $\Delta$ for an arbitrary string of quantifiers, and $t$ and $s$ for arbitrary terms of a given language. For an ordinal $\beta$ or a cardinal $\kappa$, we denote by $\beta^{+}$ and $\kappa^{+}$ their respective successors.
@@
We take the basic first-order language to be composed of individual variables, negation, conjunction, equality, and the universal quantifier (where the other connectives and quantifiers may be defined as usual). For an ordered set $\mathcal{A}$, we denote by $\mathrm{coi}(\mathcal{A})$ and $\mathrm{cof}(\mathcal{A})$ its coinitiality and cofinality, respectively. For a signature $\sigma$, we denote by $\mathcal{L}^{\sigma}$ the first-order language induced by it. As usual, we use a calligraphic letter $\mathcal{A}$ to denote a structure, and the respective italicized letter $A$ to denote its corresponding domain. Given a constant $\mathrm{c}$, function symbol $\mathrm{F}$ or relation symbol $\mathrm{R}$ of $\sigma$, and a $\sigma$-structure $\mathcal{A}$, we denote by $\mathrm{c}^{\mathcal{A}}$, $\mathrm{F}^{\mathcal{A}}$ and $\mathrm{R}^{\mathcal{A}}$ their interpretations in $\mathcal{A}$, respectively. Whenever we are working with some model, we shall state things like $\mathcal{A}\models\varphi[a_{1},...,a_{n}]$, for $a_{0},...,a_{n}\in A$, to mean $\varphi$ is a formula with at most $n$ free variables $x_{1},...,x_{n}$, and $\varphi$ is true in $\mathcal{A}$ when the $x_{i}$ are assigned the respective values $a_{i}$. Likewise, given a $\sigma$-term $t$, we shall denote by $t^{\mathcal{A}}(a_{1},...,a_{n})$ the element of $\mathcal{A}$ obtained by interpreting $t$ with each of its free variables $x_{i}$ being assigned the value $a_{i}$ – for example, in $\mathbb{N}$, $x[2]+y[3]$ is the element $2+^{\mathbb{N}}3=5$.
@@ 2. A hierarchy of cumulative powers
Let $\sigma=\langle\{F_{i}\}_{i\in I},\{R_{i}\}_{i\in J},\{c_{i}\}_{i\in K},\mathrm{ar}\rangle$ be a signature. [^4] Where $\mathrm{ar}:(\{F_{i}\}_{i\in I}\cup\{R_{i}\}_{i\in J})\to\mathbb{N}$ is the arity function. Let also $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of sets, $\mathcal{A}=\langle A,\langle\mathrm{F}_{i}\rangle_{i\in I},\langle\mathrm{R}_{i}\rangle_{i\in J},\langle\mathrm{c}_{i}\rangle_{i\in K}\rangle$ be a $\sigma$-structure, where $A$ is the underlying set, $\mathrm{F}_{i}:A^{\mathrm{ar}(F_{i})}\to A$ a $\mathrm{ar}(F_{i})$-nary function, $\mathrm{R}_{i}\subseteq A^{\mathrm{ar}(R_{i})}$ a $\mathrm{ar}(R_{i})$-ary relation, and $\mathrm{c}_{i}\in A$ the interpretation of the constant $c_{i}$. The cumulative power hierarchy generated by $\mathcal{A}$ is composed of the following structures, for each $\beta\in\mathbf{On}$:
@@
Notice the general scheme of the definition above: for a given atom $\varphi$ and elements $a_{1},...,a_{n}\in\mathcal{F}_{\beta}(A)$, $\varphi(a_{1},...,a_{n})$ is the case iff for any $j\in I_{\delta-1}$, $\varphi\big(\vartheta^{\delta}_{j}(a_{1}),...,\vartheta^{\delta}_{j}(a_{n})\big)$ holds, where $\delta=\linebreak\max\{\rho(a_{i})\}_{1\leq i\leq n}$. [^5] Although, in the case of the extended functions, one more constraint is imposed – and equality, of course, does not accord with that. Let us call such sort of definition a hereditary definition. We shall shortly see what may be said of structures defined hereditarily.
@@
    Notice that means the definition of $\mathrm{F}_{\beta}(a,b)$ when $\rho(a)\neq\rho(b)$ is such that $\mathrm{F}_{\beta}(a,b)=\mathrm{F}_{\gamma}(a,b)$, for every $\gamma\in[\delta,\beta]$. The intuitive idea in the definitions, represented by the function $\vartheta$, consists in going back to where the largest element is defined and using the operation there, which by assumption is already defined. In the definition of the relations, an equivalent definition would need to break down into the levels of each member of any $n$-tuple. For example, for a binary $\mathrm{R}$, an equivalent definition is:
@@
    We may see the amount of clauses in a definition like that grows proportionally to the number of permutations of an $n$-tuple, so that using the $\vartheta$ function allows for a more economical definition. Likewise, the definition of the set defined by the constants may be easily calculated to be:
@@
By induction on $\beta$. The base case is trivial, and for the inductive step it suffices to notice the definition of $\mathrm{F}_{\beta}$ works iff the $\mathrm{F}_{\alpha}$ are defined for all $\alpha<\beta$, so that closure is inherited pointwisely. The limit case is straightforward, as two elements in the limit case must be members of some iteration at which the operation is defined for them. ∎